OpenStudy (anonymous):

simplify 2 over x + 2 over x plus 1 - 2 over x plus 2

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

\[\frac{\frac{2}{x+2}}{x+1-\frac{2}{x+2}}\] Is this what you mean? It's hard to interpret your question without parentheses.